# A LoD of Gaussians: Unified Training and Rendering for Ultra-Large Scale Reconstruction with External Memory

### v3

@@ A.5 Evaluation Details
For OctreeGS (OctreeGS), we follow their instructions on training custom datasets. Note that OctreeGS does not perform any scene division and uses the memory-intensive ScaffoldGS (ScaffoldGS) method, making it unsuitable for scenes of this scale.
@@
We choose CityGaussian (CityGaussian) (specifically the 1.2 version of the repository) as a baseline instead of CityGaussianV2 (CityGaussianV2), as the code release for CityGaussianV2 does not yet support level-of-detail rendering. We follow the instructions on training large datasets and use the parameters and chunk split from the provided configuration file for the MatrixCity-Aerial dataset (which covers the same region as the MC-small-city+ dataset, but without the street views), as they produced better metrics than the default parameters. Note that the low number of Gaussians in the final model is due to CityGS discarding most (about 90%) of the trained Gaussians after chunk training in order to avoid artifacts. For campus and small city, we used a $4\times 4$ and $2\times 2$ chunk split to match the number of chunks as closely as possible with the 1212/44 chunks used by H-3DGS (Hierarchical).
@@
We follow the instructions of Hierarchical for running the method on large scenes. Note that we do not make use of their particular COLMAP pipeline for MC-smaller-city+, as it only works with unbroken camera paths. We disable exposure optimization for evaluation in accordance with the instructions. The partitioning step of H-3DGS requires point correspondences not provided by the dataset. To substitute, we use the camera poses provided by the dataset and generate the sparse point cloud and correspondences from scratch using COLMAP. On scenes of this scale, COLMAP output contains significant noise, which also leads to reduced quality metrics for our method.
H-3DGS achieves drastically worse PSNR and SSIM results on the full campus dataset, compared to the single-chunk results, due to a slight perspective distortion that occurs specifically on this dataset (it did not occur on small city). This can be reproduced by evaluating the trained model on their campus dataset, both of which are available on their website. As such, we include the results, but encourage visual comparison in Figure 6.
@@
We expand the configuration provided for a single chunk of the MatrixCity small-scene dataset (named ours/large_scene/block_A) with a chunk split of $3\times 4$ and $5\times 5$ for MC-smaller-city+ and MC-small-city+ (shown in Figure 22) respectively. The partitions of the MC-small-city+ scene Each chunk is trained for 60000 iterations and then fine-tuned for another 40000.
Figure 22: Partitions of HorizonGS on the MC-small-city+ scene.
@@
VastGaussian (VastGaussian) would have presented a natural comparison point, but there is currently no official code release available.
@@
We provide detailed configuration files for all of the scenes as supplemental material. In general, we perform 60k iterations of coarse training (except for 100k on MC-small-city+) and then 150k iterations of fine training (except 250k on campus and MC-smaller-city+ and 500k on MC-small-city+). We shoose SH degree 1 and half the densifaction gradient threshold as opposed to H-3DGS. We pick a cache size of 15000000 Gaussians.
A.6 Implementation Details
@@
Our general code structure is based on H-3DGS (Hierarchical). In particular, we reuse the hierarchy creator, but replace the cut procedure and chunk based training, as well as the rasterizer for the one used in 3DGS. We also use the densification implementation of 3DGSasMCMC. All of the code is implemented in PyTorch and C++/CUDA. The entire source code is included as supplemental material.
@@
The required level of detail is not only dependent on a camera’s distance $d$ to the Gaussian, but also on its focal length, which we need to account for as our datasets contain views with differing focal lengths. Therefore, we choose a base focal length $f_{b}$ and use a relative distance metric $\hat{d}$ for our HSPT cuts, which is calculated for the current camera’s focal length $f_{i}$ as $\hat{d}=\frac{f_{b}}{f_{i}}d$. E.g. for a camera with double the focal length, the cut distance should be halved to account for the larger projection footprint on the image plane.
@@
A.7 Additional Densification Details
Our densification method combines the split operation from 3DGSasMCMC with the selection method from Hierarchical. Relying purely on the densification strategy of 3DGSasMCMC–applying a noise to every Gaussian’s position and an opacity/scaling loss–becomes problematic for large scenes, where a majority of Gaussians are not visible from any one view: The result is a gradual disappearance of density in the scene or extreme ”stringing” artifacts (cf. Figure 23). However, without these losses to encourage respawning of Gaussians, the densification becomes aimless and random. We implemented another strategy that only applies the losses and noise to Gaussians that affected at least one pixel in the current iteration. While this significantly improved results, we found that the MCMC densification strategy in general performs poorly when Gaussian density is low, which is necessitated by ultra-large-scale scenes. On small scenes such as the single-chunk datasets, our modified MCMC strategy slightly outperforms the presented strategy given sufficient Gaussian budget, but it lags behind in the larger scenes.
@@ A.7 Additional Densification Details
Figure 23: Without regularization, the scale- and opacity-losses of 3DGS-as-MCMC lead to either stringing or disappearance of the entire scene.
A.8 Pseudocode
@@
A.9 Supplemental Material
As supplemental material, we include the codebase used for this project, along with the configuration files for all of our scenes. We also include configuration files for our baselines, if applicable. Further, we provide videos at the following anonymous link: https://www.youtube.com/playlist?list=PLocO6QqSKZ4YN7DyuC5gWEeVciSV6RdAk. This includes fly-throughs of the evaluated scenes and a method overview video.
A.10 Declaration of LLM usage
LLMs were used to polish the language in the paper, but did not write any sections on their own. Some code sections in the implementation were generated using LLMs.

### v4

@@ A.4. Evaluation Details
For OctreeGS (Ren et al., 2024), we follow the provided instructions on training custom datasets. We trained using the suggested hyperparameters for standard scenes and those used for the MatrixCity dataset. The latter achieved higher quality metrics, which is what we report. Note that OctreeGS does not perform any scene division and uses the memory-intensive ScaffoldGS (Lu et al., 2024) method, making it unsuitable for ultra-large scale scenes.
@@
We choose CityGaussian (Liu et al., 2024a) (specifically the 1.2 version of the repository) as a baseline instead of CityGaussianV2 (Liu et al., 2024b), as the code release for CityGaussianV2 does not yet support level-of-detail rendering. We follow the instructions on training large datasets and use the parameters and chunk split from the provided configuration file for the MatrixCity-Aerial dataset (which covers the same region as the MC-small-city+ dataset, but without the street views), as they produced better metrics than the default parameters. Note that the low number of Gaussians in the final model is due to CityGS discarding most (about 90%) of the trained Gaussians after chunk training in order to avoid artifacts. For Campus and Small City, we used the same parameters, but a $4\times 4$ and $2\times 2$ chunk split to match the number of chunks as closely as possible with the $12$ and $4$ chunks respectively used by H-3DGS (Kerbl et al., 2024) and a $3\times 3$ chunk split for the Uni10k dataset.
CLM-GS
Significant modifications were necessary to evaluate CLM-GS (Zhao et al., 2026) on our large-scale datasets. In the paper, the issue of large-scale densification is sidestepped, by starting from a dense (102 million point) ground truth point cloud and disabling densification all-together for the aerial matrix-city scene. All other evaluated scenes are sufficiently small that densification is not an issue. Running the unmodified method on large-scale datasets from sparse point clouds leads to either collapsing training or pruning of all Gaussians. Thus, we disable pruning to avoid large parts of the scene disappearing and disable opacity resets, which make the training process unstable for large-scale scenes. Experimentally we found that the high densification parameters for the 28M mode of Rubble produced the best results and used it in our evaluation. We further implemented the suggestion of Kerbl et al. (2024) to densify according to maximal instead of average screen-space gradients, but did not notice significant improvements.
As suggested by the authors for large datasets, we use the clm-offload strategy with a batch size of 4, and match the number of training iterations with our method for each experiment. We enable the option to load images from disk in order to avoid going out of memory. As of writing the CLM codebase contains no possibility of rendering test images or evaluating performance metrics, so we implemented the necessary functions ourselves.
@@
We follow the instructions of Kerbl et al. (2024) for running the method on large scenes. Note that we do not make use of their particular COLMAP pipeline for MC-smaller-city+, as it only works with unbroken camera paths. We disable exposure optimization for evaluation in accordance with the instructions. The partitioning step of H-3DGS requires point correspondences not provided by the dataset. To substitute, we use the camera poses provided by the dataset and generate the sparse point cloud and correspondences from scratch using COLMAP. On scenes of this scale, COLMAP output contains significant noise, which also leads to reduced quality metrics for our method.
H-3DGS achieves drastically worse PSNR and SSIM results on the full Campus scene, compared to the single-chunk results, due to a slight perspective distortion that occurs specifically on this dataset (it did not occur on Small City). This can be reproduced by evaluating the trained model on their Campus scene, both of which are available on their website. As such, we include the results, but encourage visual comparison in Figure 12 of the main paper.
@@
We expand the configuration provided for a single chunk of the MatrixCity small-scene dataset (named
ours/large_scene/block_A) with a chunk split of $3\times 4$ and $5\times 5$ for MC-smaller-city+ and MC-small-city+, respectively, and $2\times 2$ for Uni10k. Since the method throws an error if it is not given both a set of street- and aerial images, we limit evaluation to those two scenes. Each chunk is trained for 60000 iterations and then fine-tuned for another 40000.
The chunk partitions are shown in Figure 20; as we can see, the complexity of these partitions is vastly different. For Uni10k, the simpler chunk partitioning leads to more stable training, and, as can be seen in Table 1, improved results.
Figure 20. Partitions of HorizonGS on MC-small-city+ (left) and Uni10k (right).
@@
VastGaussian (Lin et al., 2024) would have presented a natural comparison point, but there is currently no official code release available.
OccluGaussian
As of writing, the code for (Liu et al., 2025) has not been released, so we are only able to compare to their reported results on their own dataset. Because the method is specifically tailored to indoor scenes, they never published results on any outdoor street-level datasets.
@@
In general, we perform 60k iterations of coarse training (except for 100k on MC-small-city+) and then 150k iterations of fine training (except 100k on Uni10k, 250k on Campus and MC-smaller-city+ and 500k on MC-small-city+). We choose SH degree 1 and halve the densification gradient threshold as opposed to H-3DGS. We pick a cache size of 15 million Gaussians. For each scene, we pick a maximal number of Gaussians and desired SPT volume according to the scene scale. We deactivate the use of pinned CPU memory, which would accelerate memory transfers, because it would not be available in sufficient quantities on the consumer devices we are targeting. We provide detailed configuration files containing all hyperparameters for all of the scenes as supplemental material.
A.5. Implementation Details
@@
Our general code structure is based on H-3DGS (Kerbl et al., 2024). In particular, we reuse the hierarchy creator, but replace the cut procedure and chunk based training, as well as the rasterizer for the one used in Kerbl et al. (2023). We also use the densification implementation of Kheradmand et al. (2024). All of the code is implemented in PyTorch and C++/CUDA. The entire source code is included as supplemental material.
@@
The required level of detail is not only dependent on a camera’s distance $d$ to the Gaussian, but also on its focal length, which we need to account for as our datasets contain views with differing focal lengths. Therefore, we choose a base focal length $f_{b}$ and use a relative distance metric $\hat{d}$ for our HSPT cuts, which is calculated for the current camera’s focal length $f_{i}$ as $\hat{d}=\frac{f_{b}}{f_{i}}d$. For example, given a camera with double the focal length, the cut distance should be halved to account for the larger projection footprint on the image plane.
@@
A.6. Additional Densification Details
Our densification method combines the split operation from Kheradmand et al. (2024) and the selection method from Kerbl et al. (2024) with coarse-to-fine expansion of the LoD structure. Relying purely on the densification strategy of Kheradmand et al. (2024)–applying a noise to every Gaussian’s position and an opacity/scaling loss–becomes problematic for large scenes, where a majority of Gaussians are not visible from any one view: The result is a gradual disappearance of density in the scene or extreme ”stringing” artifacts (cf. Figure 21). However, without these losses to encourage respawning of Gaussians, the densification becomes aimless and random. We implemented another strategy that only applies the losses and noise to Gaussians that affected at least one pixel in the current iteration. While this significantly improved results, we found that the MCMC densification strategy in general performs poorly when Gaussian density is low, which is necessitated by ultra-large-scale scenes. On small scenes such as the single-chunk datasets, our modified MCMC strategy slightly outperforms the presented strategy given sufficient Gaussian budget, but it lags behind in the larger scenes. Avoiding destructive opacity resets, as advocated by Kerbl et al. (2023), is particularly important in our setting, as many Gaussians will no longer conform to the hierarchical structure after opacity is restored to normal.
@@ A.6. Additional Densification Details
Figure 21. Without regularization, the scale- and opacity-losses of 3DGS-as-MCMC lead to either stringing or disappearance of the entire scene.
A.7. Pseudocode